OpenStudy (anonymous):

@hoblos : Which equation best represents the following situation? A 10 pound box of chocolates from Chocolate to Go contains a mixture of their $2.50 per pound Dark Chocolate and their 3.25 per pound cashew cluster. How many pounds of each will be in the box if it cost $16?

OpenStudy (anonymous):

2.50x + 3.25(10 - x) = 16

OpenStudy (hoblos):

well.. i think something is wrong !! if the 10 pounds are all of the cheaper chocolate the total price would be $25 !! so it can not be only $16

OpenStudy (anonymous):

That is what it says ...

OpenStudy (hoblos):

what are the choices ?

OpenStudy (anonymous):

2.50x + 3.25(10 - x) = 16 2.50(10) + 3.25x = 16 2.50x + 3.25(10) = 16 2.50(x) + 3.25(16) = 10x

OpenStudy (hoblos):

2.50x + 3.25(10 - x) = 16 represents the situation it does not make sense but it represent the situation though.. so it is the answer